# Discrete dark matter with light Dirac neutrinos

### v1

@@ Abstract
We propose a new realisation of light Dirac neutrino mass and dark matter (DM) within the framework of a non-Abelian discrete flavour symmetry $A_{4}$. In addition to $A_{4}$, we also consider a $Z_{2}$ and an unbroken global lepton number symmetry $U(1)_{L}$ to keep unwanted terms away while guaranteeing the Dirac nature of light neutrinos. The field content, their transformations and flavon vacuum alignments are chosen in such a way that the type-I Dirac seesaw generates only one light Dirac neutrino mass while the other two masses arise from scotogenic contributions at one-loop. The symmetry breaking of $A_{4}$ leaves a remnant $\mathcal{Z}_{2}$ symmetry responsible for stabilising DM. Dirac nature of light neutrinos introduces additional relativistic degrees of freedom $\Delta N_{\rm eff}$ within reach of cosmic microwave background experiments.
@@ I Introduction
Origin of light neutrino mass and mixing Zyla et al. (2020) has been one of the longstanding puzzles in particle physics. While neutrinos remain massless within the framework of the standard model (SM) of particle physics, several beyond standard model (BSM) frameworks have been proposed to explain non-zero neutrino mass and mixing. Canonical seesaw frameworks Minkowski (1977); Gell-Mann et al. (1979); Mohapatra and Senjanovic (1980); Sawada and Sugamoto (1979); Yanagida (1980); Schechter and Valle (1980); Mohapatra and Senjanovic (1981); Lazarides et al. (1981); Wetterich (1981); Schechter and Valle (1982); Foot et al. (1989) involve inclusion of heavy fields which couple to the SM leptons. These seesaw mechanisms and several of their descendants predict Majorana light neutrino masses which violate lepton number by two units. However, there has been no confirmation about the Majorana nature of light neutrinos with experiments searching for neutrinoless double beta decay (NDBD) continuing to report null results. While null results at NDBD experiments do not necessarily imply Dirac nature of light neutrinos, there has been growing interest in Dirac neutrino models of late. After a few initial attempts in this direction Roncadelli and Wyler (1983); Roy and Shanker (1984); Babu and He (1989); Peltoniemi et al. (1993), several new proposals to realise sub-eV scale Dirac neutrino masses Centelles Chuliá et al. (2017a); Aranda et al. (2014); Chen et al. (2016); Ma et al. (2015); Reig et al. (2016); Wang and Han (2016); Wang et al. (2017, 2006); Gabriel and Nandi (2007); Davidson and Logan (2009, 2010); Bonilla and Valle (2016); Farzan and Ma (2012); Bonilla et al. (2016); Ma and Popov (2017); Ma and Sarkar (2018); Borah (2016); Borah and Dasgupta (2016); Borah and Dasgupta (2017a, b); Centelles Chuliá et al. (2017b); Bonilla et al. (2018); Memenga et al. (2013); Borah and Karmakar (2018); Centelles Chuliá et al. (2018a, b); Han and Wang (2018); Borah et al. (2018); Borah and Karmakar (2019); Borah et al. (2019); Centelles Chuliá et al. (2019); Jana et al. (2019); Nanda and Borah (2019); Guo and Han (2020); Bernal et al. (2021); Borah et al. (2022a); Li et al. (2022); Dey and Roy (2024); Singh et al. (2024) have appeared recently. Most of these works, in addition to the inclusion of new fields, also consider new discrete or continuous symmetries in order to forbid unwanted terms in the Lagrangian. In the spirit of Dirac seesaw for light neutrinos, such unwanted terms include the bare mass term of Majorana type for singlet right handed neutrinos $(\nu_{R})$ and direct coupling of $\nu_{R}$ with the SM lepton and Higgs doublets.
@@
Another observed phenomena which can not be explained in the SM is the presence of dark matter (DM) in the Universe. Its presence is supported by astrophysical observations at different scales Zwicky (1933); Rubin and Ford (1970); Clowe et al. (2006) together with cosmological experiments like PLANCK, WMAP predicting around 26.8% of the present Universe to be made up of DM Zyla et al. (2020); Aghanim et al. (2020). While weakly interacting massive particle (WIMP) has been the most widely studied particle DM framework, null results at direct detection experiments Aalbers et al. (2022) have also led to growing interest in alternative scenarios where DM is more feebly coupled to the SM, like the feebly interacting massive particle (FIMP). Recent reviews of WIMP and FIMP can be found in Arcadi et al. (2017) and Bernal et al. (2017) respectively.
@@
Motivated by this, we propose a new flavour symmetric origin of light Dirac neutrino mass and dark matter in this work. We consider the popular non-Abelian discrete flavour symmetry group of $A_{4}$ augmented with an additional discrete $Z_{2}$ and global lepton number $U(1)_{L}$ symmetries. The $A_{4}$ flavour symmetry gets spontaneously broken by flavon fields while generating light Dirac neutrino mass and mixing. For a review of non-Abelian discrete flavour symmetries and their consequences, see Ref. Altarelli and Feruglio (2010); King and Luhn (2013); Petcov (2018); Chauhan et al. (2023); Ding and Valle (2024) and references therein. Unlike popular DM scenarios with additional stabilising symmetries, here we have stable DM by virtue of a remnant $\mathcal{Z}_{2}$ symmetry naturally emerging from $A_{4}$ breaking. Similar scenario for Majorana light neutrino, coined as discrete dark matter was proposed in Hirsch et al. (2010) and subsequently explored within different contexts Meloni et al. (2011); Boucenna et al. (2011); Adulpravitchai et al. (2011); Eby and Frampton (2012); Boucenna et al. (2012); Hamada et al. (2014); Lamprea and Peinado (2016); De La Vega et al. (2019); Ganguly et al. (2022); Bonilla et al. (2023); Ganguly et al. (2023); Kumar et al. (2024). We propose a realisation of this discrete dark matter paradigm with light Dirac neutrinos in a minimal $A_{4}$ flavour symmetric setup. We show that neutrino oscillation data can be satisfied after taking tree-level and one-loop contributions to light Dirac neutrino masses. While the model predicts either scalar or fermion DM with the desired relic abundance, light Dirac neutrinos lead to observable $\Delta N_{\rm eff}$ at future cosmic microwave background (CMB) experiments like CMB Stage IV (CMB-S4) Abazajian et al. (2019), SPT-3G Benson et al. (2014), Simons Observatory Ade et al. (2019) and CMB-HD Aiola et al. (2022).
@@ II The Model
$m_{D}=v_{\eta}\begin{pmatrix}y_{1}&0&0\\ y_{2}&0&0\\ y_{3}&0&0\end{pmatrix},m^{\prime}_{D}=v_{\chi}\begin{pmatrix}y^{\prime}_{1}&y^{\prime}_{2}&y^{\prime}_{3}\\ 0&0&0\\ 0&0&0\end{pmatrix},M_{R}=\begin{pmatrix}M_{N}&0&0\\ 0&M_{N}&0\\ 0&0&M_{N}\end{pmatrix};$ (4)
where the vacuum expectation values (VEVs) of the $A_{4}$ scalar triplets are considered to be $\langle\eta\rangle=(\langle\eta_{1}\rangle,\langle\eta_{2}\rangle,\langle\eta_{3}\rangle)=(\frac{v_{\eta}}{\sqrt{2}},0,0)$ and $\langle\chi\rangle=(\langle\chi_{1}\rangle,\langle\chi_{2}\rangle,\langle\chi_{3}\rangle)=(\frac{v_{\chi}}{\sqrt{2}},0,0)$. Denoting the VEV of the neutral component of the Higgs doublet $H$ as $v$, we have the constraint $\sqrt{v^{2}_{\eta}+v^{2}}=246$ GeV from the requirement of desired electroweak symmetry breaking (EWSB). Therefore, at three level, via type-I seesaw like scenario (shown in the left panel of Fig. 1) , the light Dirac neutrino mass matrix can be written as
@@
At one-loop, all the $A_{4}$ triplet scalars $\eta=(\eta_{1},\eta_{2},\eta_{3})$ and $\chi=(\chi_{1},\chi_{2},\chi_{3})$ irrespective of their VEVs can contribute to light neutrino mass, as shown in the right panel of Fig. 1. This contribution is similar to the Dirac scotogenic model discussed in several earlier works Gu and Sarkar (2008); Farzan and Ma (2012); Borah and Dasgupta (2016); Ma and Popov (2017); Borah and Dasgupta (2017b); Wang et al. (2017); Ma (2019a, b, 2020); Leite et al. (2020); Guo and Han (2020); Bernal et al. (2021); Chowdhury et al. (2022); Borah et al. (2022b, 2024a). The one-loop contribution to light Dirac neutrino mass can be estimated as
@@
with the details of the components $(m_{\nu})_{ij}$ being given in Appendix C. This mass matrix has rank 3 and predict three non-zero mass eigenvalues, in general. Also, it does not possess any specific symmetric structure and hence can be fitted to neutrino mixing data by appropriate choice of parameters, as we discuss below.
@@ III Fit to Neutrino data
data Esteban et al. (2020), the allowed ranges for $y_{i},y^{\prime}_{i}$ are given in $y^{\prime}_{1}-y_{1}$, $y^{\prime}_{2}-y_{2}$ and $y^{\prime}_{3}-y_{3}$ planes respectively, plotted in Fig. 2. In this benchmark scenario, $y^{\prime}_{i}$ are found to be in the range $0.6\times 10^{-11}\leq y^{\prime}_{1}\leq 10^{-10}$, $2.3\times 10^{-11}\leq y^{\prime}_{2}\leq 8\times 10^{-11}$, $2.35\times 10^{-11}\leq y^{\prime}_{3}\leq 8\times 10^{-11}$ and corresponding ranges for $y_{i}$ are found to be $0.038\leq y_{1}\leq 0.055$, $0.07\leq y_{2}\leq 0.1$ and $0.07\leq y_{3}\leq 0.098$ respectively. To be in agreement with neutrino oscillation data, other parameters are restricted within the range $3\,{\rm TeV}\leq M_{N}\leq 5$ TeV, $130\leq(v_{\eta}/{\rm GeV})\leq 240$, $180\leq(v_{\chi}/{\rm GeV})\leq 250$ with the loop functions defined in Eq. (7) having the magnitude $\mathcal{F}_{i}\sim\mathcal{O}(10^{-4})$. With the estimation of the parameters appearing in the neutrino mass matrix, we are now in a position to show the correlation among neutrino oscillation parameters as shown in Fig. 3. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for Case-A and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $(0.2-1.6)$ radian, $(0.075-0.089)$ eV and $(8.5-9.7)$ meV respectively.
Case-B: We now consider the benchmark scenario, $y_{i}\ll y^{\prime}_{i}$. Using similar methodology, we can again constrain the Yukawa couplings $y_{i},y^{\prime}_{i}$ imposing the constraints on neutrino masses and mixing angles Esteban et al. (2020).
@@
and $0.01\leq y^{\prime}_{1}\leq 0.1$, $0.03\leq y^{\prime}_{2}\leq 0.1$, $0.04\leq y^{\prime}_{3}\leq 0.1$ respectively. In order to satisfy the neutrino oscillation data, other parameters are restricted within the range $6\,{\rm TeV}\leq M_{N}\leq 10$ TeV, $120\leq(v_{\eta}/{\rm GeV})\leq 240$, $100\leq(v_{\chi}/{\rm GeV})\leq 700$. Corresponding loop functions $\mathcal{F}_{i}$, defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $5\times 10^{-4}\leq\mathcal{F}_{1}\leq 7\times 10^{-4}$, $3\times 10^{-4}\leq\mathcal{F}_{2}\leq 10^{-3}$, $8\times 10^{-4}\leq\mathcal{F}_{3}\leq 1.3\times 10^{-3}$. After estimating the parameters relevant for light neutrino mass matrix, we proceed to find some correlations among neutrino oscillation parameters as shown in Fig. 5. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the scenario $y_{i}\ll y^{\prime}_{i}$ and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range ±$\pm(0-1.5)$ radian, $(0.068-0.11)$ eV and $(8.4-10.5)$ meV respectively.
Case-C Finally, we now consider the third benchmark scenario where $y_{i}\simeq y^{\prime}_{i}$. In this scenario, one can also satisfy the observed constraints on neutrino masses and mixing
@@
where the associated Yukawa couplings are found to be in the range $3.56\times 10^{-7}\leq y_{1}\leq 4.1\times 10^{-7}$, $7.36\times 10^{-7}\leq y_{2}\leq 7.4\times 10^{-7}$, $7.58\times 10^{-7}\leq y_{3}\leq 7.63\times 10^{-7}$ and $6.18\times 10^{-6}\leq y^{\prime}_{1}\leq 6.3\times 10^{-6}$, $3\times 10^{-6}\leq y^{\prime}_{2}\leq 5\times 10^{-6}$ and $4.1\times 10^{-6}\leq y^{\prime}_{3}\leq 4.25\times 10^{-6}$ respectively. These allowed ranges are plotted in $y^{\prime}_{i}-y_{i}$ planes as given in Fig. 6.
@@
To fit the model with the neutrino oscillation data in this scenario, other parameters are restricted within the range $11.4\,{\rm TeV}\leq M_{N}\leq 14$ TeV, $170\leq(v_{\eta}/{\rm GeV})\leq 240$, $1\leq(v_{\chi}/{\rm TeV})\leq 1.4$. Similarly, loop functions $\mathcal{F}_{i}$ defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $10^{-3}\leq\mathcal{F}_{1}\leq 1.2\times 10^{-3}$, $4\times 10^{-4}\leq\mathcal{F}_{2}\leq 5\times 10^{-4}$, $10^{-4}\leq\mathcal{F}_{3}\leq 1.7\times 10^{-4}$. Similar to the other two cases, here also we show the correlation among neutrino oscillation parameters in Fig. 7. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the current scenario and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $-(1.57-1.5)$ radian, $(0.082-0.091)$ eV and $(8.8-9.8)$ meV respectively.

### v2

@@ Abstract
We propose a new realisation of light Dirac neutrino mass and dark matter (DM) within the framework of a non-Abelian discrete flavour symmetry based on $A_{4}$ group. In addition to $A_{4}$, we also consider a $Z_{2}$ and an unbroken global lepton number symmetry $U(1)_{L}$ to keep unwanted terms away while guaranteeing the Dirac nature of light neutrinos. The field content, their transformations and flavon vacuum alignments are chosen in such a way that the type-I Dirac seesaw generates only one light Dirac neutrino mass while the other two masses arise from scotogenic contributions at one-loop. This leads to the Dirac scoto-seesaw framework, a generalisation of the widely studied scoto-seesaw model to Dirac neutrinos. The symmetry breaking of $A_{4}$ leaves a remnant $\mathcal{Z}_{2}$ symmetry responsible for stabilising DM. Dirac nature of light neutrinos introduces additional relativistic degrees of freedom $\Delta N_{\rm eff}$ within reach of cosmic microwave background experiments.
@@ I Introduction
Origin of light neutrino mass and mixing ParticleDataGroup:2020ssz has been one of the longstanding puzzles in particle physics. While neutrinos remain massless within the framework of the standard model (SM) of particle physics, several beyond standard model (BSM) frameworks have been proposed to explain non-zero neutrino mass and mixing. Canonical seesaw frameworks Minkowski:1977sc ; GellMann:1980vs ; Mohapatra:1979ia ; Sawada:1979dis ; Yanagida:1980xy ; Schechter:1980gr ; Mohapatra:1980yp ; Lazarides:1980nt ; Wetterich:1981bx ; Schechter:1981cv ; Foot:1988aq involve inclusion of heavy fields which couple to the SM leptons. These seesaw mechanisms and several of their descendants predict Majorana light neutrino masses which violate lepton number by two units. However, there has been no confirmation about the Majorana nature of light neutrinos with experiments searching for neutrinoless double beta decay (NDBD) continuing to report null results. While null results at NDBD experiments do not necessarily imply Dirac nature of light neutrinos, there has been growing interest in Dirac neutrino models of late. After a few initial attempts in this direction Roncadelli:1983ty ; Roy:1983be ; Babu:1988yq ; Peltoniemi:1992ss , several new proposals to realise sub-eV scale Dirac neutrino masses Chulia:2016ngi ; Aranda:2013gga ; Chen:2015jta ; Ma:2015mjd ; Reig:2016ewy ; Wang:2016lve ; Wang:2017mcy ; Wang:2006jy ; Gabriel:2006ns ; Davidson:2009ha ; Davidson:2010sf ; Bonilla:2016zef ; Farzan:2012sa ; Bonilla:2016diq ; Ma:2016mwh ; Ma:2017kgb ; Borah:2016lrl ; Borah:2016zbd ; Borah:2016hqn ; Borah:2017leo ; CentellesChulia:2017koy ; Bonilla:2017ekt ; Memenga:2013vc ; Borah:2017dmk ; CentellesChulia:2018gwr ; CentellesChulia:2018bkz ; Han:2018zcn ; Borah:2018gjk ; Borah:2018nvu ; Borah:2019bdi ; CentellesChulia:2019xky ; Jana:2019mgj ; Nanda:2019nqy ; Guo:2020qin ; Bernal:2021ezl ; Borah:2022obi ; Li:2022chc ; Dey:2024ctx ; Singh:2024imk have appeared recently. Most of these works, in addition to the inclusion of new fields, also consider new discrete or continuous symmetries in order to forbid unwanted terms in the Lagrangian. In the spirit of Dirac seesaw for light neutrinos, such unwanted terms include the bare mass term of Majorana type for singlet right handed neutrinos $(\nu_{R})$ and direct coupling of $\nu_{R}$ with the SM lepton and Higgs doublets.
@@
Another observed phenomena which can not be explained in the SM is the presence of dark matter (DM) in the Universe. Its presence is supported by astrophysical observations at different scales Zwicky:1933gu ; Rubin:1970zza ; Clowe:2006eq together with cosmological experiments like PLANCK, WMAP predicting around 26.8% of the present Universe to be made up of DM ParticleDataGroup:2020ssz ; Planck:2018vyg . While weakly interacting massive particle (WIMP) has been the most widely studied particle DM framework, null results at direct detection experiments LUX-ZEPLIN:2022qhg have also led to growing interest in alternative scenarios where DM is more feebly coupled to the SM, like the feebly interacting massive particle (FIMP). Recent reviews of WIMP and FIMP can be found in Arcadi:2017kky and Bernal:2017kxu respectively.
@@
Motivated by this, we propose a new flavour symmetric origin of light Dirac neutrino mass and dark matter in this work. We consider the popular non-Abelian discrete flavour symmetry group of $A_{4}$ augmented with an additional discrete $Z_{2}$ and global lepton number $U(1)_{L}$ symmetries. The $A_{4}$ flavour symmetry gets spontaneously broken by flavon fields while generating light Dirac neutrino mass and mixing. For a review of non-Abelian discrete flavour symmetries and their consequences, see Ref. Altarelli:2010gt ; King:2013eh ; Petcov:2017ggy ; Chauhan:2023faf ; Ding:2024ozt and references therein. Unlike popular DM scenarios with additional stabilising symmetries, here we have stable DM by virtue of a remnant $\mathcal{Z}_{2}$ symmetry naturally emerging from $A_{4}$ breaking. Similar scenario for Majorana light neutrino, coined as discrete dark matter was proposed in Hirsch:2010ru and subsequently explored within different contexts Meloni:2010sk ; Boucenna:2011tj ; Adulpravitchai:2011ei ; Eby:2011qa ; Boucenna:2012qb ; Hamada:2014xha ; Lamprea:2016egz ; DeLaVega:2018bkp ; Bonilla:2023pna ; Kumar:2024zfb . We propose a realisation of this discrete dark matter paradigm with light Dirac neutrinos in a minimal $A_{4}$ flavour symmetric setup. We show that neutrino oscillation data can be satisfied after taking tree-level and one-loop contributions to light Dirac neutrino masses. Interestingly, this also leads to the generalisation of the popular Scoto-Seesaw scenario Rojas:2018wym to Dirac neutrinos, and can be termed as Dirac Scoto-Seesaw. While the model predicts either scalar or fermion DM with the desired relic abundance, light Dirac neutrinos lead to observable $\Delta N_{\rm eff}$ at future cosmic microwave background (CMB) experiments like CMB Stage IV (CMB-S4) Abazajian:2019eic , SPT-3G SPT-3G:2014dbx , Simons Observatory SimonsObservatory:2018koc and CMB-HD CMB-HD:2022bsz .
@@ II The Model
$m_{D}=\frac{v_{\eta}}{\sqrt{2}}\begin{pmatrix}y_{1}&0&0\\ y_{2}&0&0\\ y_{3}&0&0\end{pmatrix},m^{\prime}_{D}=\frac{v_{\chi}}{\sqrt{2}}\begin{pmatrix}y^{\prime}_{1}&y^{\prime}_{2}&y^{\prime}_{3}\\ 0&0&0\\ 0&0&0\end{pmatrix},M_{R}=\begin{pmatrix}M_{N}&0&0\\ 0&M_{N}&0\\ 0&0&M_{N}\end{pmatrix};$ (4)
respectively, where the vacuum expectation values (VEVs) of the $A_{4}$ scalar triplets are considered to be $\langle\eta\rangle=(\langle\eta_{1}\rangle,\langle\eta_{2}\rangle,\langle\eta_{3}\rangle)=(\frac{v_{\eta}}{\sqrt{2}},0,0)$ and $\langle\chi\rangle=(\langle\chi_{1}\rangle,\langle\chi_{2}\rangle,\langle\chi_{3}\rangle)=(\frac{v_{\chi}}{\sqrt{2}},0,0)$. Denoting the VEV of the neutral component of the Higgs doublet $H$ as $v$, we have the constraint $\sqrt{v^{2}_{\eta}+v^{2}}=246$ GeV from the requirement of desired electroweak symmetry breaking (EWSB). Therefore, at three level, via type-I seesaw like scenario (shown in the left panel of Fig. 1) , the light Dirac neutrino mass matrix can be written as
@@
At one-loop, all the $A_{4}$ triplet scalars $\eta=(\eta_{1},\eta_{2},\eta_{3})$ and $\chi=(\chi_{1},\chi_{2},\chi_{3})$ irrespective of their VEVs can contribute to light neutrino mass, as shown in the right panel of Fig. 1. This contribution is similar to the Dirac scotogenic [^1] Scotogenic model was originally proposed in Tao:1996vb ; Ma:2006fn for Majorana neutrinos. model discussed in several earlier works Gu:2007ug ; Farzan:2012sa ; Borah:2016zbd ; Ma:2016mwh ; Borah:2017leo ; Wang:2017mcy ; Ma:2019iwj ; Ma:2019yfo ; Ma:2019coj ; Leite:2020wjl ; Guo:2020qin ; Bernal:2021ezl ; Chowdhury:2022jde ; Borah:2022phw ; Borah:2022enh . The one-loop contribution to light Dirac neutrino mass can be estimated as
@@
with the details of the components $(m_{\nu})_{ij}$ being given in Appendix C. For simplicity, the Yukawa couplings appearing in the neutrino mass matrix are considered to be real. Thus, CP violation is introduced only through the complex parameter $\omega$, appearing here as a consequence of the considered $A_{4}$ discrete flavour symmetry. This mass matrix has rank 3 and predict three non-zero mass eigenvalues, in general. Also, it does not possess any specific symmetric structure and hence can be fitted to neutrino mixing data by appropriate choice of parameters, as we discuss below. Interestingly, in the limit $y_{2}=y_{3}$ and $y^{\prime}_{2}=y^{\prime}_{3}$, the effective mass matrix given in Eq. (11) yields $\mu-\tau$ reflection symmetric mass matrix Harrison:2002et predicting fixed values for $\theta_{23}(=\pi/4)$ and $\delta_{\rm CP}(=\pm\pi/2$), see Appendix C for details. Since type-I seesaw generates only one mass scale with the other being generated by the scotogenic contribution, our framework provides a generalisation of the scoto-seesaw mechanism Rojas:2018wym ; Mandal:2021yph ; Barreiros:2020gxu ; Barreiros:2022aqu ; Ganguly:2022qxj ; Ganguly:2023jml ; Kumar:2023moh to light Dirac neutrinos. Similar to earlier works on Majorana neutrinos, Dirac scoto-seesaw framework can explain the two mass scales in neutrino sector namely, atmospheric and solar mass scales, from tree level (type-I seesaw) and radiative (scotogenic) contributions respectively.
@@ III Fit to Neutrino data
data Esteban:2020cvm , the allowed ranges for $y_{i},y^{\prime}_{i}$ are given in $y^{\prime}_{1}-y_{1}$, $y^{\prime}_{2}-y_{2}$ and $y^{\prime}_{3}-y_{3}$ planes respectively, plotted in Fig. 2. In this benchmark scenario, $y^{\prime}_{i}$ are found to be in the range $6\times 10^{-11}\leq y^{\prime}_{1}\leq 10^{-10}$, $3\times 10^{-11}\leq y^{\prime}_{2,3}\leq 8.5\times 10^{-11}$ and corresponding ranges for $y_{i}$ are found to be $0.03\leq y_{1}\leq 0.05$, $0.052\leq y_{2}\leq 0.1$ and $0.05\leq y_{3}\leq 0.1$ respectively. To be in agreement with neutrino oscillation data, other parameters are restricted within the range $2.8\,\leq(M_{N}/{\rm TeV})\leq 5$, $20\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.22\leq(v_{\chi}/{\rm TeV})\leq 1.8$ with the loop functions defined in Eq. (7) having the magnitude $\mathcal{F}_{i}\sim\mathcal{O}(10^{-4})$. With the estimation of the parameters appearing in the neutrino mass matrix, we are now in a position to show the correlation among neutrino oscillation parameters as shown in Fig. 3. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for Case-A and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $(0.43-1.6)$ radian, $(0.06-0.09)$ eV and $(8.4-9.7)$ meV respectively.
Case-B: We now consider the benchmark scenario, $y_{i}\ll y^{\prime}_{i}$. Using similar methodology, we can again constrain the Yukawa couplings $y_{i},y^{\prime}_{i}$ imposing the constraints on neutrino masses and mixing angles Esteban:2020cvm .
@@
and $0.01\leq y^{\prime}_{1}\leq 0.1$, $0.03\leq y^{\prime}_{2}\leq 0.1$, $0.04\leq y^{\prime}_{3}\leq 0.1$ respectively. In order to satisfy the neutrino oscillation data, other parameters are restricted within the range $6\,\leq(M_{N}/{\rm TeV})\leq 10$, $20\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.2\leq(v_{\chi}/{\rm TeV})\leq 1.8$. Corresponding loop functions $\mathcal{F}_{i}$, defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $5\times 10^{-4}\leq\mathcal{F}_{1}\leq 7\times 10^{-4}$, $3\times 10^{-4}\leq\mathcal{F}_{2}\leq 10^{-3}$, $8\times 10^{-4}\leq\mathcal{F}_{3}\leq 1.3\times 10^{-3}$. After estimating the parameters relevant for light neutrino mass matrix, we proceed to find some correlations among neutrino oscillation parameters as shown in Fig. 5. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the scenario $y_{i}\ll y^{\prime}_{i}$ and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range ±$\pm(0-1.6)$ radian, $(0.068-0.11)$ eV and $(8.4-10.5)$ meV respectively.
Case-C: Finally, we now consider the third benchmark scenario where $y_{i}\simeq y^{\prime}_{i}$. In this scenario, one can also satisfy the observed constraints on neutrino masses and mixing
@@
where the associated Yukawa couplings are found to be in the range $3\times 10^{-7}\leq y_{1}\leq 4.6\times 10^{-7}$, $6.9\times 10^{-7}\leq y_{2}\leq 7.9\times 10^{-7}$, $7.1\times 10^{-7}\leq y_{3}\leq 8.1\times 10^{-7}$ and $5.6\times 10^{-6}\leq y^{\prime}_{1}\leq 6.8\times 10^{-6}$, $2.9\times 10^{-6}\leq y^{\prime}_{2}\leq 5.5\times 10^{-6}$ and $3.6\times 10^{-6}\leq y^{\prime}_{3}\leq 4.75\times 10^{-6}$ respectively. These allowed ranges are plotted in $y^{\prime}_{i}-y_{i}$ planes as given in Fig. 6.
@@
To fit the model with the neutrino oscillation data in this scenario, other parameters are restricted within the range $8\,\leq(M_{N}/{\rm TeV})\leq 16$, $35\leq(v_{\eta}/{\rm GeV})\leq 240$, $0.23\leq(v_{\chi}/{\rm TeV})\leq 1.8$. Similarly, loop functions $\mathcal{F}_{i}$ defined in Eq. (7), appearing in the Dirac scotogenic contribution are found to be in the range $10^{-3}\leq\mathcal{F}_{1}\leq 1.2\times 10^{-3}$, $4\times 10^{-4}\leq\mathcal{F}_{2}\leq 7\times 10^{-4}$, $10^{-4}\leq\mathcal{F}_{3}\leq 1.7\times 10^{-4}$. Similar to the other two cases, here also we show the correlation among neutrino oscillation parameters in Fig. 7. In this figure, we have plotted the correlations for $\delta_{\rm CP}-\sin^{2}\theta_{23}$, $\sum m_{i}-m_{1}$, $\sum m_{i}-\sin^{2}\theta_{23}$ planes for the current scenario and predicted ranges for $\delta_{CP}$, $\sum m_{i}$ and $m_{1}$ are found to be in the range $(1.2-1.58)$ and $-(1-1.58)$ radian, $(0.082-0.11)$ eV and $(8.8-11)$ meV respectively.